OpenStudy (anonymous):

kia drew triangle pqr and triangle stu so that angle p is congruent to angle s, q is congruent to t, pr=12 and su=3. if pqr and stu are similar identify the similar postulate that applies A cannot be determined B similar side angle side C similar angle angle D similar side side side

OpenStudy (anonymous):

is the answer a

OpenStudy (anonymous):

can you please explain to me why did u use triangle, im really confuse of it sorry. :D

OpenStudy (anonymous):

on my quiz they have the triangle shape next to pqr and stu

OpenStudy (anonymous):

ohhh ok thnks infamous21 :D

OpenStudy (anonymous):

i'll try that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if two pairs of angles are congruent, the third pair of angles will be also congruent, right?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

The sum of angles of a triangle is 180. p =s, q=t, so 180-(p+q) = 180 -(s+t), that is, angle r = angle u

OpenStudy (anonymous):

ok so its d sss

OpenStudy (anonymous):

not quite. you don't have data on more than one side. You have information about two angles.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

it was c